# An Early Look at the Performance of IGRINS–2 at Gemini-North with Application to the ultrahot Jupiter, WASP-33 b

### v1

@@ 5 Summary
We successfully detected clear cross-correlation signals from these molecular species in the dayside atmosphere of WASP-33 b with a combined SNR of 7.5. The detection was achieved by cross-correlating the post-SVD residual spectra with atmospheric models convolved to match the instrumental resolution of R=45,000. The results confirmed the presence of molecular species that were consistent with previous results from other high-resolution spectroscopy studies (e.g., Nugroho et al., 2021; Cont et al., 2021, 2022a, 2022b; Yan et al., 2022; Finnerty et al., 2023), supporting the conclusion of a thermally inverted atmosphere for WASP-33 b. The detected planetary trail, which aligns with the expected radial velocity curve, further corroborates the planetary origin of the signal and underscores the precision of IGRINS-2 in tracking Doppler-shifted atmospheric features.
@@
Our results are in good agreement with previous studies on ultrahot Jupiters (e.g., Finnerty et al., 2023; Brogi et al., 2023), further supporting the reliability of atmospheric models for these extreme environments. In particular, our detection of CO and H2O is consistent with theoretical predictions for high-temperature atmospheres where strong molecular excitation occurs (Madhusudhan et al., 2011). The presence of OH in the atmosphere adds to the growing body of evidence for complex chemical processes at play in these planetary systems, particularly in atmospheres experiencing significant photodissociation and thermal inversion.
@@ Acknowledgments
Y.H.C. thanks Stefan Pelletier, Woowon Byun, and Jaerim Koo for improving the manuscript. This work used the Immersion Grating Infrared Spectrograph 2 (IGRINS-2) developed and built by a collaboration between Korea Astronomy and Space Science Institute (KASI) and the International Gemini Observatory. Based on observations obtained at the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation on behalf of the Gemini Observatory partnership: the U.S. National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). This work was enabled by observations made from the Gemini North telescope, located within the Maunakea Science Reserve and adjacent to the summit of Maunakea. We are grateful for the privilege of observing the Universe from a place that is unique in both its astronomical quality and its cultural significance.

### v3

@@ 5 Summary
We successfully detected clear cross-correlation signals from these molecular species in the dayside atmosphere of WASP-33 b with a combined SNR of 7.4. The detection was achieved by cross-correlating the post-SVD residual spectra with atmospheric models convolved to match the instrumental resolution of R=45,000. The results confirmed the presence of molecular species that were consistent with previous results from other high-resolution spectroscopy studies (e.g., Nugroho et al., 2021; Cont et al., 2021, 2022a, 2022b; Yan et al., 2022; Finnerty et al., 2023), supporting the conclusion of a thermally inverted atmosphere for WASP-33 b. The detected planetary trail, which aligns with the expected radial velocity curve, further corroborates the planetary origin of the signal and underscores the precision of IGRINS-2 in tracking Doppler-shifted atmospheric features.
@@
Our results are in good agreement with previous studies on ultrahot Jupiters (e.g., Finnerty et al., 2023; Brogi et al., 2023), further supporting the reliability of atmospheric models for these extreme environments. In particular, our detection of CO and H2O is consistent with theoretical predictions for high-temperature atmospheres where strong molecular excitation occurs (Madhusudhan et al., 2011). The clear detection of OH additionally provides indirect evidence of active thermal dissociation processes, as OH is a product of H2O dissociation under intense stellar irradiation (Brogi et al., 2023). This further highlights the chemically complex and dynamic nature of ultra-hot Jupiter atmospheres.
@@ Acknowledgments
Y.H.C. thanks Stefan Pelletier, Woowon Byun, and Jaerim Koo for improving the manuscript. This research was partly supported by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2025-1-830-05, Project No. 2025-1-860-02, Project No. 2025-1-868-02) supervised by the Korea AeroSpace Administration. This work used the Immersion Grating Infrared Spectrograph 2 (IGRINS-2) developed and built by a collaboration between Korea Astronomy and Space Science Institute (KASI) and the International Gemini Observatory. Based on observations obtained at the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation on behalf of the Gemini Observatory partnership: the U.S. National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). This work was enabled by observations made from the Gemini North telescope, located within the Mauna Kea Science Reserve and adjacent to the summit of Mauna Kea. We are grateful for the privilege of observing the Universe from a place that is unique in both its astronomical quality and its cultural significance.